# OC4-ReID: Occluded Cloth-Changing Person Re-Identification

### v3

@@ Abstract
The study of Cloth-Changing Person Re-identification (CC-ReID) focuses on retrieving specific pedestrians when their clothing has changed, typically under the assumption that the entire pedestrian images are visible. Pedestrian images in real-world scenarios, however, are often partially obscured by obstacles, presenting a significant challenge to existing CC-ReID systems. In this paper, we introduce a more challenging task termed Occluded Cloth-Changing Person Re-Identification (OC4-ReID), which simultaneously addresses two challenges of clothing changes and occlusion. Concretely, we construct two new datasets, Occ-LTCC and Occ-PRCC, based on original CC-ReID datasets to include random occlusions of key pedestrians components (e.g., head, torso). Moreover, a novel benchmark is proposed for OC4-ReID incorporating a Train-Test Micro Granularity Screening (T2MGS) module to mitigate the influence of occlusion and proposing a Part-Robust Triplet (PRT) loss for partial features learning. Comprehensive experiments on the proposed datasets, as well as on two CC-ReID benchmark datasets demonstrate the superior performance of proposed method against other state-of-the-art methods. The codes and datasets are available at: https://github.com/1024AILab/OC4-ReID.
@@ I Introduction
Given the significant impact of occlusion on CC-ReID, we introduce a novel and more challenging task, termed Occluded Cloth-Changing Person Re-identification (OC4-ReID). This task aims to identify pedestrians while overcoming the dual chanllenges of occlusion and clothing changes.
Figure 1: We construct OC4-ReID dataset based on the CC-ReID dataset by introducing random occlusion on various human body parts (e.g., head, arms) . Fair training and testing results indicate that traditional CC-ReID methods (e.g., CAL [9] and AIM [10]) perform worse on the OC4-ReID dataset.
@@
In this paper, we propose two specialized OC4-ReID datasets and a simple yet effective benchmark for OC4-ReID task. Firstly, We adopt a pre-trained human parser [11] to locate the body components of pedestrians in PRCC [12] and LTCC [13] datasets, then randomly select one component to occlude, constructing the OC4-ReID datasets. Secondly, to mitigate the impact of occlusion and mine discriminative local representations, we propose a Train-Test Micro Granularity Screening (T2MGS) module, in which a quality predictor is maintained to evaluate each micro granularity features. During testing, low-quality features are discarded to suppress occlusion. Thirdly, we also propose a Part-Robust Triplet loss (PRT) to supervise partial features, which aims to combine all the body parts and mine truly discriminative representations. In addition, PRT is robust to clothing features in pedestrians’ partial appearances and emphasises the extraction of clothing-unrelated identity information. The main contributions of this paper are summarized as follows:
@@ II OC4-ReID Dataset
As shown in Fig. 2, to facilitate the study of OC4-ReID, we customise two datasets Occ-PRCC and Occ-LTCC based on existing cloth-changing datasets [12, 13]. Motivated by prior works [14, 15], we use black patches to represent occlusions. Specifically, a pre-trained human parser, SCHP [11], is employed to locate the key components of pedestrians (i.e., head, torso, upper or lower arms, upper or lower legs). We then randomly select a component to get the occlusion target $f_{\text{mask}}$, and this process is detailed as $f_{\text{mask}}=\mathcal{F}_{\text{rand}}\left(g_{\theta}(x)\right)$, where $x$ denotes the input images, $g_{\theta}(\cdot)$ is the processes of SCHP inference, and $\mathcal{F}_{\text{rand}}(\cdot)$ refers to the random selection process. Note that occlusions in real-world do not exactly match the segmentation regions, therefore we perform a pooling and an up-sampling operation on the segmentation results to obtain more arbitrary occlusion shapes. Finally, we combine the occlusions and original images to get the processed images $f_{\text{occ}}$:
@@
where $\mathcal{F}_{\text{pool}}(\cdot)$ and $\mathcal{F}_{\text{up}}(\cdot)$ denotes the average pooling and up-sampling operations, and $\mathcal{F}_{\text{fusion}}(\cdot,\cdot)$ represents the images fusion process.
@@ III-A Train-Test Micro Granularity Screening Module
Let $F\in\mathbb{R}^{C\times H\times W}$ denotes the output features of the backbone, where $C$, $H$, and $W$ are the number of channels, height, and width. We divide $F$ into $k$ horizontal partitions denoted as $f_{1},f_{2},\dots,f_{k}\in\mathbb{R}^{C\times\frac{H}{k}\times W}$. For each partition, we maintain a quality predictor, which comprise two convolutional layers, a batch normalization layer, and a sigmoid activation layer. The above process can be formulated as:
@@
During the testing phase, for the quality prediction results $f_{i}^{\phi}(i\in\{1,2,\dots,k\})$ of $k$ partitions, we conduct a micro granularity screening in the channel dimension. Specifically, channels with quality scores lower than the threshold $\lambda$ (set to 0.35 in this paper) are discarded. Through this, our model can adaptively discard occluded features while focusing on discriminative representations. The detailed procedure is outlined in Algorithm 1.
@@ III-B Part-Robust Triplet Loss
In part-based ReID methods [18, 19, 20, 21], identity-related loss is directly used to supervise partial features. However, this approach encounters limitations in the OC4-ReID scenario: (i) Generally, pedestrian clothing is more likely to act as a discriminative feature and promote the ReID task. However, clothing features are interference information in OC4-ReID, supervising the clothing region with identity-related loss impairs model performance. (ii) Occluded regions mainly contain blurred and low-quality features with little identity information. Using identity-related loss for supervision introduces misleading information. To address these issues, we propose the Part-Robust Triplet (PRT) Loss, which considers the average distances of all partial features and focuses on mining the overall discriminative representations:
@@
where $\mathcal{F}_{\text{max}}(\cdot,\cdot)$ returns the maximum value, and $M$ is the margin in triplet loss, which is set to 0.5 in this paper. For global features $f_{g}$ and $f_{g}^{w}$, we utilize cross-entropy loss [17] as the identity loss to supervise, which are denoted as $\mathcal{L}_{ID}^{g}$ and $\mathcal{L}_{ID}^{p}$. Additionally, similar to [9], we introduce a clothing-based adversarial loss $\mathcal{L}_{CA}$ that collaborates with the clothing classification loss $\mathcal{L}_{C}$ [17] to learn the clothing-unrelated features. The final total loss is calculated as follows:
@@ IV-B Implementation Details
We implement our method with PyTorch and train models on two NVIDIA RTX 4090 with 24 Gigabyte memory. ResNet-50 [11] pre-trained on ImageNet is selected as the backbone. All the input images are resized to $384\times 192$. Random horizontal flipping, cropping and erasing are utilized for data augmentation. The batch size is set to 64, and initial learning rate is set to $3.5\times 10^{-4}$, which decays at 30th and 50th epochs with a decay factor of 0.1 (120 epochs in total).
@@ IV-C Compare with State-of-the-art Methods
Result on Occ-PRCC and Occ-LTCC: As shown in Table I, our model outperforms existing state-of-the-art methods on the OC4-ReID datasets. Taking the Occ-PRCC dataset as an example, our model achieves 58.81% Rank-1 and 56.52% mAP, which is 11.71% and 11.54% higher than the second-best method. Similarly, we also achieve advanced performance on the Occ-LTCC dataset, with Rank-1 and mAP improving by 9.22% and 8.46% compared to the second-best method. The experimental results demonstrate the superior performance of our method in handling the OC4-ReID task, because the following two reasons: (i) By conducting screening mechanisms on local and channel features, T2MGS effectively suppresses low-quality features while further mining fine-grained discriminative representations. (ii) Using PRT loss to supervise partial features, this key design motivates the model to focus on truly discriminative features at each training step, which in turn mitigates the effect of occlusion and clothing features.
@@ IV-D Ablation Study of The Proposed Method
TABLE II: Ablation Experiments of Our Method on PRCC Dataset.
@@
Table II reports the ablation experimental results of our method in Occ-PRCC. Our baseline achieve 45.34% Rank@1 and 43.21% mAP accuracy. When T2MGS is introduced for occlusion mitigation (Index 1), the Rank@1 and mAP are improved by 9.39% and 11.44%, respectively. Then, we use classic triplet loss [16] for supervising partial features (Index 2), which reduce the Rank@1 and mAP by 2.61% and 2.54%. The results show that acting triplet loss directly on partial features inevitably damages the feature extraction process and introduces interference information. Finally, when replacing triplet loss with the proposed PRT loss (Index 3), our method achieves optimal performance with Rank@1 and mAP of 58.81% and 56.52%.
@@ IV-E Sensitivity Analysis of Hyper-Parameter Setting
To explore the optimal value of threshold $\lambda$ and partition number $k$, we conduct parameter-sensitivity experiments on the Occ-PRCC dataset. We first set $\lambda$ and $k$ to 0.15 and 2, then increase them by 0.05 and 2, respectively. As shown in Fig. 4 (a), the model’s performance reduces when the threshold $\lambda$ once exceeds 0.35, which indicates that visible features are removed along with occluded features. In addition, as illustrated in Fig. 4 (b), the model achieve best performance when the number of partitions $k$ is set to 6.
@@ V Conclusion
In this paper, we propose the task of cloth-changing person re-identification under occlusion conditions, termed OC4-ReID, for the first time. To support this new task, we construct two specialized datasets, Occ-PRCC and Occ-LTCC. To address the challenges presented in this task, we design a new benchmark, which includes a Train-Test Micro Granularity Screening (T2MGS) module. This module effectivelymitigates occlusion while facilitating the extraction of fine-grained cloth-unrelated features. Furthermore, we propose Part-Robust Triplet (PRT) loss, which supervises the partial features learning properly. Experimental results on both CC-ReID and OC4-ReID datasets demonstrate the effectiveness of our proposed method.

### v4

@@ Abstract
The study of Cloth-Changing Person Re-identification (CC-ReID) focuses on retrieving specific pedestrians when their clothing has changed, typically under the assumption that the entire pedestrian images are visible. Pedestrian images in real-world scenarios, however, are often partially obscured by obstacles, presenting a significant challenge to existing CC-ReID systems. In this paper, we introduce a more challenging task termed Occluded Cloth-Changing Person Re-Identification (OC4-ReID), which simultaneously addresses two challenges of clothing changes and occlusions. Concretely, we construct two new datasets, Occ-LTCC and Occ-PRCC, based on original CC-ReID datasets to include random occlusions of key pedestrians components (e.g., head, torso). Moreover, a novel benchmark is proposed for OC4-ReID incorporating a Train-Test Micro Granularity Screening (T2MGS) module to mitigate the influence of occlusions and proposing a Part-Robust Triplet (PRT) loss for partial features learning. Comprehensive experiments on the proposed datasets, as well as on two CC-ReID benchmark datasets demonstrate the superior performance of proposed method against other state-of-the-art methods. The codes and datasets are available at: https://github.com/1024AILab/OC4-ReID.
@@ I Introduction
Given the significant impact of occlusions on CC-ReID, we introduce a novel and more challenging task, termed Occluded Cloth-Changing Person Re-identification (OC4-ReID). This task aims to identify pedestrians while overcoming the dual challenges of occlusions and clothing changes.
Figure 1: We construct OC4-ReID dataset based on the CC-ReID dataset by introducing random occlusions on various human body parts (e.g., head, arms) . Fair training and testing results indicate that traditional CC-ReID methods (e.g., CAL [9] and AIM [10]) perform worse on the OC4-ReID dataset.
@@
In this paper, we propose two specialized OC4-ReID datasets and a simple yet effective benchmark for OC4-ReID task. Firstly, We adopt a pre-trained human parser [11] to locate the body components of pedestrians in PRCC [12] and LTCC [13] datasets, then randomly select one component to occlude, constructing the OC4-ReID datasets. Secondly, to mitigate the impact of occlusion and mine discriminative local representations, we propose a Train-Test Micro Granularity Screening (T2MGS) module, in which a quality predictor is maintained to evaluate each micro granularity features. During testing, low-quality features are discarded to suppress occlusions. Thirdly, we also propose a Part-Robust Triplet loss (PRT) to supervise partial features, which aims to combine all the body parts and mine truly discriminative representations. In addition, PRT is robust to clothing features in pedestrians’ partial appearances and emphasises the extraction of clothing-unrelated identity information. The main contributions of this paper are summarized as follows:
@@ II OC4-ReID Dataset
As shown in Fig. 2, to facilitate the study of OC4-ReID, we customise two datasets Occ-PRCC and Occ-LTCC based on existing cloth-changing datasets [12, 13]. Motivated by prior works [14, 15], we use black patches to represent occlusions. Specifically, a pre-trained human parser, SCHP [11], is employed to locate the key components of pedestrians (i.e., head, torso, upper or lower arms, upper or lower legs). We then randomly select a component to get the occlusion target $f_{\text{mask}}$, and this process is detailed as $f_{\text{mask}}=\mathcal{F}_{\text{rand}}\left(g_{\theta}(x)\right)$, where $x$ denotes the input image, $g_{\theta}(\cdot)$ is the processes of SCHP inference, and $\mathcal{F}_{\text{rand}}(\cdot)$ refers to the random selection process. Note that occlusions in real-world do not exactly match the segmentation regions, therefore we perform a pooling and an up-sampling operation on the segmentation results to obtain more arbitrary occlusion shapes. Finally, we combine the occluded and original image to get the processed image $f_{\text{occ}}$:
@@
where $\mathcal{F}_{\text{av-pool}}(\cdot)$ and $\mathcal{F}_{\text{up}}(\cdot)$ denotes the average pooling and up-sampling operations, and $\mathcal{F}_{\text{fusion}}(\cdot,\cdot)$ represents the images fusion process.
@@ III-A Train-Test Micro Granularity Screening Module
Let $F\in\mathbb{R}^{C\times H\times W}$ denotes the output feature of the backbone, where $C$, $H$, and $W$ are the number of channels, height, and width. We divide $F$ into $k$ horizontal partitions denoted as $f_{1},f_{2},\dots,f_{k}\in\mathbb{R}^{C\times\frac{H}{k}\times W}$. For each partition, we maintain a quality predictor, which comprise two convolutional layers, a batch normalization layer, and a sigmoid activation layer. The above process can be formulated as:
@@
During the testing phase, for the quality prediction result $f_{i}^{\phi}(i\in\{1,2,\dots,k\})$ of $k$ partitions, we conduct a micro granularity screening in the channel dimension. Specifically, channels with quality scores lower than the threshold $\lambda$ (set to 0.35 in this paper) are discarded. Through this, our model can adaptively discard occluded features while focusing on discriminative representations. The detailed procedure is outlined in Algorithm 1.
@@ III-B Part-Robust Triplet Loss
In part-based ReID methods [18, 19, 20, 21], identity-related triplet loss [16] is directly used to supervise partial features. However, these approaches encounter limitations in the OC4-ReID scenario: (i) Generally, pedestrian clothing is more likely to act as a discriminative feature and promote the ReID task. However, clothing features are interference information in OC4-ReID, supervising the clothing region with identity-related loss impairs model performance. (ii) Occluded regions mainly contain blurred and low-quality features with little identity information. Using identity-related loss for supervision introduces misleading information. To address these issues, we propose the Part-Robust Triplet (PRT) Loss, which considers the average distances of all partial features and focuses on mining the overall discriminative representations:
@@
where $\mathcal{F}_{\text{max}}(\cdot,\cdot)$ returns the maximum value, and $M$ is the margin in triplet loss, which is set to 0.3 in this paper. For global features $f_{g}$ and $f_{g}^{w}$, we utilize cross-entropy loss [17] as the identity loss to supervise, which are denoted as $\mathcal{L}_{ID}^{g}$ and $\mathcal{L}_{ID}^{p}$. Additionally, similar to [9], we introduce a clothing-based adversarial loss $\mathcal{L}_{CA}$ that collaborates with the clothing classification loss $\mathcal{L}_{C}$ [17] to learn the clothing-unrelated features. The final total loss is calculated as follows:
@@ IV-B Implementation Details
We implement our method with PyTorch and train the model on two NVIDIA RTX 4090 with 24 Gigabyte memory. ResNet-50 [11] pre-trained on ImageNet is selected as the backbone. All the input images are resized to $384\times 192$. Random horizontal flipping, cropping and erasing are utilized for data augmentation. The batch size is set to 64, and initial learning rate is set to $3.5\times 10^{-4}$, which decays at 30th and 50th epochs with a decay factor of 0.1 (120 epochs in total).
@@ IV-C Compare with State-of-the-art Methods
Result on Occ-PRCC and Occ-LTCC: As shown in Table I, our model outperforms existing state-of-the-art methods on the OC4-ReID datasets. Taking the Occ-PRCC dataset as an example, our model achieves 58.81% Rank-1 and 56.52% mAP, which is 11.71% and 11.54% higher than the second-best method. Similarly, we also achieve advanced performance on the Occ-LTCC dataset, with Rank-1 and mAP improving by 9.22% and 8.46% compared to the second-best method. The experimental results demonstrate the superior performance of our method in handling the OC4-ReID task, because the following two reasons: (i) By conducting screening mechanisms on local and channel features, T2MGS effectively suppresses low-quality features while further mining fine-grained discriminative parts. (ii) Using PRT loss to supervise partial features, this key design motivates the model to focus on truly discriminative features at each training step, which in turn mitigates the effect of occlusions and clothing features.
@@ IV-D Ablation Study of The Proposed Method
TABLE II: Ablation Experiments of Our Method on Occ-PRCC Dataset.
@@
Table II reports the ablation experimental results of our method in Occ-PRCC. Our baseline achieves 45.34% Rank@1 and 43.21% mAP accuracy. When T2MGS is introduced for occlusion mitigation (Index 1), the Rank@1 and mAP are improved by 9.39% and 11.44%, respectively. Then, we use classic triplet loss [16] for supervising partial features (Index 2), which reduce the Rank@1 and mAP by 2.61% and 2.54%. The results show that acting triplet loss directly on partial features inevitably damages the feature extraction process and introduces interference information. Finally, when replacing triplet loss with the proposed PRT loss (Index 3), our method achieves optimal performance with Rank@1 and mAP of 58.81% and 56.52%.
@@ IV-E Sensitivity Analysis of Hyper-Parameter Setting
To explore the optimal value of threshold $\lambda$ and partition number $k$, we conduct parameter-sensitivity experiments on the Occ-PRCC dataset. We first set $\lambda$ and $k$ to 0.15 and 2, then increase them by 0.05 and 2, respectively. As shown in Fig. 4 (a), the model’s performance reduces when the threshold $\lambda$ once exceeds 0.35, which indicates that visible features are removed along with occluded features. In addition, as illustrated in Fig. 4 (b), the model achieves best performance when the number of partitions $k$ is set to 6.
@@ V Conclusion
In this paper, we propose the task of cloth-changing person re-identification under occlusion conditions, termed OC4-ReID, for the first time. To support this new task, we construct two specialized datasets, Occ-PRCC and Occ-LTCC. To address the challenges presented in this task, we design a new benchmark, which includes a Train-Test Micro Granularity Screening (T2MGS) module. This module effectively mitigates occlusions while facilitating the extraction of fine-grained cloth-unrelated features. Furthermore, we propose a Part-Robust Triplet (PRT) loss, which supervises the partial features learning properly. Experimental results on both CC-ReID and OC4-ReID datasets demonstrate the effectiveness of our proposed method.